OpenStudy (blake57roger):

Find the exact area of the circle with the given radius. r = 4.5 m

OpenStudy (ipwnbunnies):

Area = pi*r^2

OpenStudy (ipwnbunnies):

For an exact answer, leave pi as pi.

OpenStudy (anonymous):

\[\text{Area}=\pi \left(\frac{45}{10}\right)^2 \text{meters}^2 \]

OpenStudy (blake57roger):

whats the anwser

OpenStudy (anonymous):

A=pir^2, where 4.5 is your r >_>

OpenStudy (mathmale):

@blake57roger : It's up to you to do the math indicated. Since you want the "exact" answer, do not use an approximation for Pi; instead, keep Pi in your answer. Evaluate 45^2 and 10^2 and then divide the 2nd into the first. Or simply leave the result given you by robtobey exactly as it is (as it involves NO rounding off).